i need to see exactly how to foil (3+2h*0.839)^2=33
Answer by stanbon(75887)

i need to see exactly how to solve
(3+2h*0.839)^2=33
------
3+2h*0.839 = sqrt(33)
-----
1.678h = sqrt(33)-3
---
h = (sqrt(33)+3)/1.678
----
h = 5.211
==============
Cheers,
Stan H.